OpenStudy (anonymous):

Brandon needs $480 to buy a TV and stereo system for his room. He received $60 in cash for birthday presents. He plans to save $30 per week from his part-time job. To find how many weeks w it will take to have $480, solve 60 + 30w = 480.

OpenStudy (anonymous):

wait.......whts ur question

OpenStudy (anonymous):

14 weeks man.

OpenStudy (anonymous):

60 + 30w =480 -60 -60 30w = 420 w = 14